Paul Company had 100,000 shares of common stock outstanding on January 1,2018. On September 30, 2018, Paul sold 58,000 shares of common stock for cash. Paul also had 15,000 shares of convertible preferred stock outstanding throughout 2018. The preferred stock is $100 par, 7%, and is convertible into 3 shares of common for each share of preferred. Paul also had 600, 996, convertible bonds outstanding throughout 2018. Each $1,000 bond is convertible into 30 shares of common stock. The bonds sold originally at face value. Reported net income for 2018 was $400,000 with a 40% tax rate. Common shareholders received $3.00 per share dividends after preferred dividends were paid in 2018. Required Compute basic and diluted earnings per share for 2018. (Round your answers to 2 decimal places.) share Basic Diluted
Answer #1

Weighted common share = 100000+(58000*3/12) = 114500 Shares

a) Basic earning per share = (Net income-Preferred dividend) / Share outstanding

                                        = (400000-105000)/114500

Basic earning per share = 2.58 per share

b) Diluted earning per share = Adjusted net income/Adjusted diluted share

Adjusted net income = 400000+(600*1000*9%*60%) = 432400

Adjusted diluted shares = 114500+(15000*3)+(600000/30) = 179500

Diluted earning per share = 432400/179500 = 2.41 per share